A 0.5kg cart is connected to light spring for which the force constant is 20n/m on a horizontal frictionless air track . Calculate the total energy of the system and the maximum speed of the cart if the amplitude of the motion is 3.0cm
1
Expert's answer
2021-09-24T09:25:47-0400

The total energy of the system


"E=kx^2/2=20\cdot0.03^2/2=0.009\ (J)"


The maximum speed of the cart

"E=mv^2/2\to v=\sqrt{2E/m}=\sqrt{2\cdot0.009/0.5}=0.19\ (m/s)=19\ (cm/s)"